Thermochemical changes: Standard Enthalpies of Formation


Fe2O3(s)+3CO(g)→3CO2(g)+2Fe(s) Standard enthalpy change of this reaction is: -24.8kJ


Question: what is the mass of carbon dioxide produced in the above reaction when 250 kJ of energy are released?


1
Expert's answer
2020-10-26T14:49:05-0400

Standard molar enthalpy of this reaction is: ΔrHmº = -24.8 kJ/3mol

The amount of moles produced when the amount of energy is released:

n = 250 kJ/(-24.8 kJ/3mol) = 30.24193548 mol

The molar mass for each mole of substance:

M (CO2) = (1(12.01)) + (2(16.00)) = 44.01 g/mol

30.24193548 mol x 44.01 g/mol = 1330.94758 g

1330.94758 g/1000 g = 1.33 kg